OpenStudy (anonymous):

A coin is tossed and then a die is rolled. Find the probability of getting a 5 on the die given that the coin landed tails up.

OpenStudy (anonymous):

p(t) = 1/2 p(5) = 1/6 p(t and 5) = 1/2 x/1/6 = 1/12